# BLANKET: Anonymizing Faces in Infant Video Recordings

### v1

@@ I Introduction
Sharing datasets of human subjects, including video footage, is crucial for advancing psychological science [adolph2017video]. However, for ethical reasons, this cannot be done without anonymizing the videos. To study development, daily spontaneous recordings of infants will be needed [adolph2011sampling] to uncover developmental trajectories at multiple and nested time scales. The issues related to privacy and ethics are even more pressing when it comes to infants and children.
@@
Figure 1: Infant face anonymization. Frame of the original video (a), anonymization by the proposed BLANKET method (b), and by DeepPrivacy2 [deepprivacy2] (c). The proposed method alters the identity of the infant without introducing obvious perceptual artifacts, while keeping all other facial attributes intact, e.g. face orientation, gaze, expression. The method can handle face occlusion, which is challenging for the competing method. Original video courtesy of Stephen Julia, Max Family Fun.
@@
The challenge is that the subjects are infants or newborns. This demographic group is underrepresented in computer vision datasets and models typically struggle to generalize to unseen data. Infants have different facial features and proportions than adults. For instance, popular generative models for face synthesis, e.g. StyleGAN2 [Karras2019stylegan2] has difficulties to generate infants, since it was trained on the FFHQ dataset containing mostly photos of adult faces downloaded from Flicker. Therefore, we use, Stable Diffusion [stable-diffusion], a large scale generative model to do the inpainting in the first stage. This model was trained on a LAION dataset of billion image scale which certainly includes infants, and so produces an image with a new undistorted identity.
The new identity is inserted into every frame of the video by FaceFusion [Facefusion], which ensures efficient, seamless and temporally consistent face swap, with high-fidelity expression transfer from the original recording. See Fig. 1 for an example of anonymization performed by our method and by a recent popular method DeepPrivacy2 [deepprivacy2], a GAN-based model.
@@ II Related Work
There exists a wide variety of anonymization techniques to conceal the identity of a person in images or video recordings. We refer to a survey [overview] or a recent journal article [kim2025diffface] for a more detailed taxonomy of the methods. There are two approaches: Face obscuring and Face modifying methods.
@@ II-B Face Modifying Methods
Before deep learning, the methods utilized face-swap by finding a matching face of a different identity in a large dataset. The new face replaced the original face and was finally blended into the original image. Examples of these methods are the k-Same algorithm [k-same] and AnonySwap [anonyswap].
More recent methods often employ deep generative models for anonymization. The methods are typically based on inpainting [sun2018natural], or face swapping [natsume2019fsnet]. The generative models used are based on Generative Adversarial Networks (GANs) [wu2019privacy, maximov2020ciagan, deepprivacy, deepprivacy2], or more recently on Diffusion Models [zhao2023diffswap, kim2025diffface].
Consistent preservation of the above-mentioned facial properties is usually not guaranteed. Moreover, all the above methods work for images and not for a video recording. Unpleasant flickering artifacts would appear if used naively frame-by-frame on a video. The problem of video anonymization seems to be unexplored, although there are exceptions [balaji2021temporally]. A recent paper [rosberg2023fiva] reports promising results in videos. However, the code is not publicly available, so prospects on infant face anonymization remain unclear.
@@ III Method
Figure 2: Flowchart of the proposed video anonymization. First, an image replacing the original identity with a new compatible random identity from the first frame is created. The new face is generated by inpainting using Stable Diffusion [stable-diffusion]. Then, the new identity is swapped in every frame of the video. We propose to use FaceFusion [Facefusion], as it provides temporally consistent results while preserving original facial expressions.
Here we propose BLANKET (Baby-face Landmark-preserving ANonymization with Keypoint dEtection consisTency). The fundamental idea of this anonymization method is to employ a temporally consistent face swapping algorithm. The input consists of an original video and an image representing the new identity; see Fig. 2. This technique ensures that the original identity is seamlessly replaced with the new identity in the output video, avoiding spatial artifacts (e.g., obvious seams, lighting inconsistencies, or spurious details) and temporal inconsistencies (e.g., blinking or flickering). In addition, except for the change of identity, all other attributes of the face are preserved, including head orientation, gaze, eyeblinks, facial expressions, and emotions. The output anonymized video looks natural for a human and is usable for downstream tasks such as facial analysis and body pose estimation, in the same way as the original video. We propose using FaceFusion [Facefusion], which is a recent framework for advanced face swapping. FaceFusion detects, tracks, and aligns faces and performs face swapping. The tool incorporates lip-syncing, expression-matching, and a face enhancement model that together ensure consistency in pose, expression, and lighting.
The challenge is to find an image of a new identity that is significantly different from the original but still compatible. For example, it is not possible to perform a faceswap with the face of an adult onto a newborn. The new face should be of the same race and gender, with approximately similar geometric proportions. Therefore, inspired by [hadera-2024], we propose generating the new identity randomly using inpainting with the diffusion model, namely Stable Diffusion [stable-diffusion].
@@
The face detector provides positions of the faces as rectangular bounding boxes. We use YOLO11 [yolov11] to detect faces [yolo-face]. The face bounding boxes are used to find the locations of facial landmarks. The landmarks correspond to important facial features, such as the eyes, mouth, and nose. We use the SPIGA model [spiga] with “wflw” weights, which detects 98 distinct landmarks and predicts a head pose. Then we construct a binary mask, as a convex hull of the landmarks. The mask delineates the inner area of the face to be modified by inpainting.
Image inpainting, as described in [inpainting], is the process of filling in missing parts of an image with such content that an unknowing observer would not notice the modification. Since the inpainting has no or very limited access to the region inside the mask, the inpainting algorithm generates a random identity that matches the area outside of the mask. In this way, an anonymized image is finally generated.
We employ the generative model called Stable Diffusion [stable-diffusion] as our image generator. For generating images, we use two different checkpoints: “Realistic Vision V2.0” [^1] https://huggingface.co/SG161222/Realistic_Vision_V2.0 and “Realistic Vision V6.0 B1” [^2] https://civitai.com/models/4201/realistic-vision-v60-b1.
@@
We use a positive prompt “a face of a baby” and a standard negative prompt. We used CFG [CFG-ho] and Control-NET [Control-NET] to fit the landmarks and face geometry and empirically tuned all parameters to balance the above trade-off. The complete parameter setting can be found in the code repository.
@@ IV Experiments
In the experiments, we extensively evaluate the proposed anonymization method and compare it with a recent popular tool, DeepPrivacy2 [deepprivacy2].
@@
For evaluation, we used a subset of Infant Pose Estimation dataset by Chambers et al. [YTdataset] presented in [chambers2020computer]. Each of 46 videos of infants is approximately 100 seconds long, the total footage time is 78 minutes comprising about 140k frames.
@@ IV-A Anonymization statistics
To evaluate the level of de-identification, we use standard face recognition model, ArcFace [ArcFace]. The model outputs $L_{2}$-normalized descriptors for each face image. Small cosine distance indicates faces of the same identity, while high cosine distance occurs for different identities. Cosine distance $1-\mathbf{v_{o}}^{T}\mathbf{v_{a}}$ between pairs of frames of the original $\mathbf{v_{o}}$ and anonymized $\mathbf{v_{a}}$ identity descriptors is calculated. Tab. I shows average cosine distances calculated over frames of all test videos. It is seen that the difference between the models is not very significant, which we hypothesize is because the ArcFace model was not trained and does not generalize well to infants and newborns. So to assess the level of de-identification, we refer to our user study and qualitative results.
@@
To assess how much the anonymization methods preserve facial atributes of the original video, we measure several statistics. All facial attributes are estimated by the DeepFace library [deepface]. Gender (male, female), race (Indian, Asian, Latino Hispanic, Black, Middle Eastern, White) and emotions (sad, angry, surprise, fear, happy, disgust, neutral) are estimated in both original and anonymized videos. Tab. II shows the ratio of frames in which the classification is the same. It is seen that the proposed method preserves these attributes better, especially for emotions, due to the fidelity of the expression transfer from the original to the anonymized video.
@@
We further measure preservation of gaze, openness of eyes and mouth, see Tab. III. Gaze is a 2D vector that spans the interval $1,1]\times[-1,1]$ for the left-right and up-down directions. The gaze difference is calculated as the Eucledian vector difference between the original and anonymized video. The openness of the eyes and mouth is calculated as the aspect ratio $d_{tb}/d_{lr}$ between the top-bottom and left-right distances between the respective landmarks in the shape outline, following [soukupova2016eye]. The difference between the original and anonymized frames is calculated and averaged over all frames in the dataset. Our method achieves better preservation of the openness of the eye and mouth.
@@ IV-B Impact of anonymization to pose estimation
We evaluate performance degradation in a downstream task: Human Pose Estimation (HPE). State-of-the-art HPE methods follow a top-down pipeline, where a person detector is applied first, followed by a pose estimator for each detected bounding box. We use RTMDet-l [RTMDet] for person detection due to its efficiency and competitive accuracy. For pose estimation, we adopt the current SOTA ViTPose-b [ViTPose], trained jointly on COCO [COCO], MPII [MPII], and AIC [AIC].
@@
Tab. VI reports performance drop in detection AP for different anonymization methods. The no-anonymization baseline refers to RTMDet applied to the original images. The AP does not reach 100 due to pose-based NMS in pseudo GT generation. RTMDet has only bouning box-based NMS and its AP reaches 98.1 compared to pseudo GT. Both black rectangle and DeepPrivacy2 reduce detection quality. These methods deform the head region, often truncating bounding boxes around the head (Fig. 4). The black rectangle always alters the head shape, but detection can still succeed if ears or hair are visible. DeepPrivacy2 replaces the face but can fail completely, leading to missed detections, as shown in Fig. 4. BLANKET achieves over 90 AP compared to the original detections. The drop is caused by a different number of detections (both FPs and FNs) compared to the pseudo ground truth (Fig. 5). Note that 90 AP is a very high result – the current SOTA [coDETR] on the COCO dataset is 66.0 AP.
@@
Figure 4: Anonymization could change output of a detector. Original image (a), black rectangle anonymization (b), DeepPrivacy2 (c) and BLANKET (d) with their respective bboxes detected by RTMDet [RTMDet]. Missing face information truncates the head (b) or enlarge detection (c).
Figure 5: Deformed face could lead to wrong number of detections. Original image (a) with it’s bounding box detected by RTMDet [RTMDet], black rectangle anonymization (b), DeepPrivacy2 (c) and BLANKET (d). Missing face information in (b) and (c) causes false positive detections (purple bboxes).
Figure 6: Anonymization methods hinder facial pose estimation. Original image (a), black rectangle anonymization (b), DeepPrivacy2 (c) and BLANKET (d) with their respective poses detected by ViTpose [ViTPose]. Methods (b) and (c) dramatically hinder pose estimation on facial keypoints.
Figure 7: Face is crucial for human pose estimation in uncommon poses. Original image (a), black rectangle anonymization (b), DeepPrivacy2 (c) and BLANKET (d) with their respective poses detected by ViTpose [ViTPose]. Missing face information in (b) and (c) causes complete failure in uncommon poses.
@@ IV-C User study
We conducted a user study to assess the level of de-identification, and level of perceptual artifacts. Both of these respects are hard to measure automatically. The standard face recognition engine [ArcFace] is not reliable for infants. Perceptual artifacts, disturbing for human observers, are difficult to predict in the images automatically. We tried [zhang2023perceptual], but it did not work.
@@ V Conclusion, Discussion and Future Work
A limitation of our method is that users occasionally perceived the de-identification as insufficient, with the anonymized face resembling the original. This may stem from the low resolution of infant faces, where detailed features require zooming. Additionally, the method alters only the facial interior, retaining landmarks that inherently carry identity cues. Modifying the full facial structure is complex. An alternative is to generate identities maximally dissimilar to the original, though difficult to automate, especially since, as we verified, face recognition engines perform poorly on infants. Future work will explore metrics for selecting dissimilar yet compatible identities with minimal artifacts. Following [Subrtova-2024], we will also examine methods for detecting and correcting perceptual artifacts.

### v2

@@ I Introduction
Sharing datasets of human subjects, including video footage, is crucial for advancing psychological science [2]. However, for ethical reasons, this cannot be done without anonymizing the videos. To study development, daily spontaneous recordings of infants will be needed [1] to uncover developmental trajectories at multiple and nested time scales. The issues related to privacy and ethics are even more pressing when it comes to infants and children.
@@
Figure 1: Infant face anonymization. Frame of the original video (a), anonymization by the proposed BLANKET method (b), and by DeepPrivacy2 [12] (c). The proposed method alters the identity of the infant without introducing obvious perceptual artifacts, while keeping all other facial attributes intact, e.g. face orientation, gaze, expression. The method can handle face occlusion, which is challenging for the competing method. Original video courtesy of Stephen Julia, Max Family Fun.
@@
The challenge is that the subjects are infants or newborns. This demographic group is underrepresented in computer vision datasets and models typically struggle to generalize to unseen data. Infants have different facial features and proportions than adults. For instance, popular generative models for face synthesis, e.g. StyleGAN2 [16] has difficulties to generate infants, since it was trained on the FFHQ dataset containing mostly photos of adult faces downloaded from Flicker. Therefore, we use, Stable Diffusion [26], a large scale generative model to do the inpainting in the first stage. This model was trained on a LAION dataset of billion image scale which certainly includes infants, and so produces an image with a new undistorted identity.
The new identity is inserted into every frame of the video by FaceFusion [9], which ensures efficient, seamless and temporally consistent face swap, with high-fidelity expression transfer from the original recording. See Fig. 1 for an example of anonymization performed by our method and by a recent popular method DeepPrivacy2 [12], a GAN-based model.
@@ II Related Work
There exists a wide variety of anonymization techniques to conceal the identity of a person in images or video recordings. We refer to a survey [25] or a recent journal article [17] for a more detailed taxonomy of the methods. There are two approaches: Face obscuring and Face modifying methods.
@@ II-B Face Modifying Methods
Before deep learning, the methods utilized face-swap by finding a matching face of a different identity in a large dataset. The new face replaced the original face and was finally blended into the original image. Examples of these methods are the k-Same algorithm [23] and AnonySwap [21].
More recent methods often employ deep generative models for anonymization. The methods are typically based on inpainting [31], or face swapping [22]. The generative models used are based on Generative Adversarial Networks (GANs) [33, 20, 13, 12], or more recently on Diffusion Models [37, 17].
Consistent preservation of the above-mentioned facial properties is usually not guaranteed. Moreover, all the above methods work for images and not for a video recording. Unpleasant flickering artifacts would appear if used naively frame-by-frame on a video. The problem of video anonymization seems to be unexplored, although there are exceptions [4]. A recent paper [27] reports promising results in videos. However, the code is not publicly available, so prospects on infant face anonymization remain unclear.
@@ III Method
Figure 2: Flowchart of the proposed video anonymization. First, an image replacing the original identity with a new compatible random identity from the first frame is created. The new face is generated by inpainting using Stable Diffusion [26]. Then, the new identity is swapped in every frame of the video. We propose to use FaceFusion [9], as it provides temporally consistent results while preserving original facial expressions.
Here we propose BLANKET (Baby-face Landmark-preserving ANonymization with Keypoint dEtection consisTency). The fundamental idea of this anonymization method is to employ a temporally consistent face swapping algorithm. The input consists of an original video and an image representing the new identity; see Fig. 2. This technique ensures that the original identity is seamlessly replaced with the new identity in the output video, avoiding spatial artifacts (e.g., obvious seams, lighting inconsistencies, or spurious details) and temporal inconsistencies (e.g., blinking or flickering). In addition, except for the change of identity, all other attributes of the face are preserved, including head orientation, gaze, eyeblinks, facial expressions, and emotions. The output anonymized video looks natural for a human and is usable for downstream tasks such as facial analysis and body pose estimation, in the same way as the original video. We propose using FaceFusion [9], which is a recent framework for advanced face swapping. FaceFusion detects, tracks, and aligns faces and performs face swapping. The tool incorporates lip-syncing, expression-matching, and a face enhancement model that together ensure consistency in pose, expression, and lighting.
The challenge is to find an image of a new identity that is significantly different from the original but still compatible. For example, it is not possible to perform a faceswap with the face of an adult onto a newborn. The new face should be of the same race and gender, with approximately similar geometric proportions. Therefore, inspired by [10], we propose generating the new identity randomly using inpainting with the diffusion model, namely Stable Diffusion [26].
@@
The face detector provides positions of the faces as rectangular bounding boxes. We use YOLO11 [14] to detect faces [15]. The face bounding boxes are used to find the locations of facial landmarks. The landmarks correspond to important facial features, such as the eyes, mouth, and nose. We use the SPIGA model [24] with “wflw” weights, which detects 98 distinct landmarks and predicts a head pose. Then we construct a binary mask, as a convex hull of the landmarks. The mask delineates the inner area of the face to be modified by inpainting.
Image inpainting, as described in [5], is the process of filling in missing parts of an image with such content that an unknowing observer would not notice the modification. Since the inpainting has no or very limited access to the region inside the mask, the inpainting algorithm generates a random identity that matches the area outside of the mask. In this way, an anonymized image is finally generated.
We employ the generative model called Stable Diffusion [26] as our image generator. For generating images, we use two different checkpoints: “Realistic Vision V2.0” [^1] https://huggingface.co/SG161222/Realistic_Vision_V2.0 and “Realistic Vision V6.0 B1” [^2] https://civitai.com/models/4201/realistic-vision-v60-b1.
@@
We use a positive prompt “a face of a baby” and a standard negative prompt. We used CFG [11] and Control-NET [36] to fit the landmarks and face geometry and empirically tuned all parameters to balance the above trade-off. The complete parameter setting can be found in the code repository.
@@ IV Experiments
In the experiments, we extensively evaluate the proposed anonymization method and compare it with a recent popular tool, DeepPrivacy2 [12].
@@
For evaluation, we used a subset of Infant Pose Estimation dataset by Chambers et al. [7] presented in [6]. Each of 46 videos of infants is approximately 100 seconds long, the total footage time is 78 minutes comprising about 140k frames.
@@ IV-A Anonymization statistics
To evaluate the level of de-identification, we use standard face recognition model, ArcFace [8]. The model outputs $L_{2}$-normalized descriptors for each face image. Small cosine distance indicates faces of the same identity, while high cosine distance occurs for different identities. Cosine distance $1-\mathbf{v_{o}}^{T}\mathbf{v_{a}}$ between pairs of frames of the original $\mathbf{v_{o}}$ and anonymized $\mathbf{v_{a}}$ identity descriptors is calculated. Tab. I shows average cosine distances calculated over frames of all test videos. It is seen that the difference between the models is not very significant, which we hypothesize is because the ArcFace model was not trained and does not generalize well to infants and newborns. So to assess the level of de-identification, we refer to our user study and qualitative results.
@@
To assess how much the anonymization methods preserve facial atributes of the original video, we measure several statistics. All facial attributes are estimated by the DeepFace library [28]. Gender (male, female), race (Indian, Asian, Latino Hispanic, Black, Middle Eastern, White) and emotions (sad, angry, surprise, fear, happy, disgust, neutral) are estimated in both original and anonymized videos. Tab. II shows the ratio of frames in which the classification is the same. It is seen that the proposed method preserves these attributes better, especially for emotions, due to the fidelity of the expression transfer from the original to the anonymized video.
@@
We further measure preservation of gaze, openness of eyes and mouth, see Tab. III. Gaze is a 2D vector that spans the interval $1,1]\times[-1,1]$ for the left-right and up-down directions. The gaze difference is calculated as the Eucledian vector difference between the original and anonymized video. The openness of the eyes and mouth is calculated as the aspect ratio $d_{tb}/d_{lr}$ between the top-bottom and left-right distances between the respective landmarks in the shape outline, following [29]. The difference between the original and anonymized frames is calculated and averaged over all frames in the dataset. Our method achieves better preservation of the openness of the eye and mouth.
@@ IV-B Impact of anonymization to pose estimation
We evaluate performance degradation in a downstream task: Human Pose Estimation (HPE). State-of-the-art HPE methods follow a top-down pipeline, where a person detector is applied first, followed by a pose estimator for each detected bounding box. We use RTMDet-l [19] for person detection due to its efficiency and competitive accuracy. For pose estimation, we adopt the current SOTA ViTPose-b [34], trained jointly on COCO [18], MPII [3], and AIC [32].
@@
Tab. VI reports performance drop in detection AP for different anonymization methods. The no-anonymization baseline refers to RTMDet applied to the original images. The AP does not reach 100 due to pose-based NMS in pseudo GT generation. RTMDet has only bouning box-based NMS and its AP reaches 98.1 compared to pseudo GT. Both black rectangle and DeepPrivacy2 reduce detection quality. These methods deform the head region, often truncating bounding boxes around the head (Fig. 4). The black rectangle always alters the head shape, but detection can still succeed if ears or hair are visible. DeepPrivacy2 replaces the face but can fail completely, leading to missed detections, as shown in Fig. 4. BLANKET achieves over 90 AP compared to the original detections. The drop is caused by a different number of detections (both FPs and FNs) compared to the pseudo ground truth (Fig. 5). Note that 90 AP is a very high result – the current SOTA [38] on the COCO dataset is 66.0 AP.
@@
Figure 4: Anonymization could change output of a detector. Original image (a), black rectangle anonymization (b), DeepPrivacy2 (c) and BLANKET (d) with their respective bboxes detected by RTMDet [19]. Missing face information truncates the head (b) or enlarge detection (c).
Figure 5: Deformed face could lead to wrong number of detections. Original image (a) with it’s bounding box detected by RTMDet [19], black rectangle anonymization (b), DeepPrivacy2 (c) and BLANKET (d). Missing face information in (b) and (c) causes false positive detections (purple bboxes).
Figure 6: Anonymization methods hinder facial pose estimation. Original image (a), black rectangle anonymization (b), DeepPrivacy2 (c) and BLANKET (d) with their respective poses detected by ViTpose [34]. Methods (b) and (c) dramatically hinder pose estimation on facial keypoints.
Figure 7: Face is crucial for human pose estimation in uncommon poses. Original image (a), black rectangle anonymization (b), DeepPrivacy2 (c) and BLANKET (d) with their respective poses detected by ViTpose [34]. Missing face information in (b) and (c) causes complete failure in uncommon poses.
@@ IV-C User study
We conducted a user study to assess the level of de-identification, and level of perceptual artifacts. Both of these respects are hard to measure automatically. The standard face recognition engine [8] is not reliable for infants. Perceptual artifacts, disturbing for human observers, are difficult to predict in the images automatically. We tried [35], but it did not work.
@@ V Conclusion, Discussion and Future Work
A limitation of our method is that users occasionally perceived the de-identification as insufficient, with the anonymized face resembling the original. This may stem from the low resolution of infant faces, where detailed features require zooming. Additionally, the method alters only the facial interior, retaining landmarks that inherently carry identity cues. Modifying the full facial structure is complex. An alternative is to generate identities maximally dissimilar to the original, though difficult to automate, especially since, as we verified, face recognition engines perform poorly on infants. Future work will explore metrics for selecting dissimilar yet compatible identities with minimal artifacts. Following [30], we will also examine methods for detecting and correcting perceptual artifacts.